OpenStudy (anonymous):

write y - 9 = -(x+2) in slope-intercept form.

OpenStudy (anonymous):

y=-x+7

OpenStudy (istim):

What's slope intercept form? (y2-Y1)=m(X1-Y2) or y=mx+b?

OpenStudy (anonymous):

y=mx+b the one u mentioned is point slope form

OpenStudy (istim):

Ok, so then simplify the right side first.

OpenStudy (istim):

Then, add 9 to both sides.

OpenStudy (anonymous):

yup

OpenStudy (istim):

You good?`